# CheXmask: a large-scale dataset of anatomical segmentation masks for multi-center chest x-ray images

### v3

@@ Methods
The CheXmask dataset is structured as one file of comma-separated values (CVS) for each CXR dataset. The CSV content is explained in Table 1. We do not include images or metadata from the original datasets. We include an image ID in the first column of the CSV files, which has the same column name as the ID column of the respective dataset, and allows to match the rows in this CSV with the original dataset’s instances. We also include the pre-processed version of the masks, allowing to use all datasets in the same image resolutions (please also note that ChestX-ray8 masks are already in the desired resolution). This database [29] is available at PhysioNet [30], in the following repository: https://physionet.org/content/chexmask-cxr-segmentation-data/.
@@
Image ID
Contains references to the original images as per the original metadata, thus the column name changes across dataset.
Dice RCA (Max)
Provides the maximum Dice Similarity Coefficient for the Reverse Classification Accuracy (RCA), indicating the quality of the segmentation.
Dice RCA (Mean)
Provides the mean Dice Similarity Coefficient for the Reverse Classification Accuracy (RCA), indicating the quality of the segmentation.
Landmarks
Includes a set of points representing the contour of the organs, as obtained by the HybridGNet.
Left Lung
Contains segmentation masks of the left lung in run-length encoding (RLE).
Right Lung
Contains segmentation masks of the right lung in RLE.
Heart
Contains segmentation masks of the heart in RLE.
Height
Height of the segmentation mask, necessary to decode RLE.
Width
Width of the segmentation mask, necessary to decode RLE.
@@
As part of the technical validation on the quality of lungs and heart masks, we report a statistical description of the distribution of mean RCA-estimated DSC for all images in each dataset (Table 3). The full histograms are presented in Figure 7. For all datasets, the mean RCA-estimated DSC is higher than 0.88. As explained above, the true DSC distribution probably has a larger mean, since the RCA-estimated DSC was found to under-estimate the performance.
@@ Acknowledgements
This work was supported by Argentina’s National Scientific and Technical Research Council (CONICET), which covered the salaries of E.F., D.M. and N.G. The authors gratefully acknowledge NVIDIA Corporation with the donation of the GPUs used for this research and ANPCyT (PICT-PRH-2019-00009).

### v4

@@ Methods
The CheXmask dataset is structured as one file of comma-separated values (CSV) for each CXR dataset. The CSV content is explained in Table 1. We do not include images or metadata from the original datasets. We include an image ID in the first column of the CSV files, which has the same column name as the ID column of the respective dataset, and allows to match the rows in this CSV with the original dataset’s instances. We also include the pre-processed version of the masks, allowing to use all datasets in the same image resolutions (please also note that ChestX-ray8 masks are already in the desired resolution). This database [29] is available at PhysioNet [30], in the following repository: https://physionet.org/content/chexmask-cxr-segmentation-data/.
@@
Image ID Contains references to the original images as per the original metadata, thus the column name changes across dataset.
Dice RCA (Max) Provides the maximum Dice Similarity Coefficient for the Reverse Classification Accuracy (RCA), indicating the quality of the segmentation.
Dice RCA (Mean) Provides the mean Dice Similarity Coefficient for the Reverse Classification Accuracy (RCA), indicating the quality of the segmentation.
Landmarks Includes a set of points representing the contour of the organs, as obtained by the HybridGNet.
Left Lung Contains segmentation masks of the left lung in run-length encoding (RLE).
Right Lung Contains segmentation masks of the right lung in RLE.
Heart Contains segmentation masks of the heart in RLE.
Height Height of the segmentation mask, necessary to decode RLE.
Width Width of the segmentation mask, necessary to decode RLE.
@@
As part of the technical validation on the quality of lungs and heart masks, we report a statistical description of the distribution of mean RCA-estimated DSC for all images in each dataset (Table 3). The full histograms are presented in Figure 7. For all datasets, the mean RCA-estimated DSC is higher than 0.83. As explained above, the true DSC distribution probably has a larger mean, since the RCA-estimated DSC was found to under-estimate the performance.
@@ Acknowledgements
This work was supported by Argentina’s National Scientific and Technical Research Council (CONICET), which covered the salaries of E.F., D.M., L.M. and N.G. The authors gratefully acknowledge NVIDIA Corporation with the donation of the GPU computing used for this research and Agencia Nacional de Promoción de la Investigación, el Desarrollo Tecnológico y la Innovación (Agencia I+D+i PICT).